# PDNNet: PDN-Aware GNN-CNN Heterogeneous Network for Dynamic IR Drop Prediction

### v1

@@ IV-A Experimental Setting
TABLE III: The statistics of Mini-CircuitNet. Samples randomly sample from the origin superset. Note that training & testing datasets have no intersection in designs.
{tabu}
l—c—c—ccc Design Number of Samples Netlist Statistics
#Cells #Nets Cell Area ($\mu m^{2}$)
Train RISCY-a 20 44,836 80,287 65,739
RISCY-b 22 30,207 58,452 69,779
RISCY-FPU-a 21 61,677 106,429 75,985
RISCY-FPU-b 17 47,130 84,676 80,030
Test zero-riscy-a 13 35,017 67,472 58,631
zero-riscy-b 7 20,350 45,599 62,648
Datasets. CircuitNet [33, 35] is the large-scale public dataset in development that provides IC design used in the real industrial scenario. Unlike most previous work using small datasets for validation, CircuitNet-N28 (28nm version) provides more than 10,000 samples from 6 basic RTL designs (54 different synthesized netlists) with a wide variety of macros (3 types), frequencies (50/200/500 MHz), and back-end flow (variant in settings of utilization, macro placement, power mesh, filler insertion, etc), which is favorably appropriate to examine the actual performance of dynamic IR drop prediction models. We follow the same protocol as CircuitNet [33], to split the training and testing sets. That is, 4 designs (RISCY-a, RISCY-b, RISCY-FPU-a, RISCY-FPU-b) with 7078 samples are for training, and the rest 2 designs (zero-riscy-a and zero-riscy-b) with 3164 samples are for testing. CircuitNet-N28 leverages Innovus and Voltus flow to obtain vectorless dynamic IR drop solutions. In our experiment, we extract the PDN features from the Design Exchange Format (DEF) files while incorporating the power information that was initially provided to complete our PDNGraph construction. Besides, to avoid unnecessary training burden, we also collect a subset of CircuitNet, named Mini-CircuitNet, for conducting the ablation studies in Section IV-C. To do so, we sample 100 total samples from the original CircuitNet dataset, where 80 and 20 samples are for the training and testing phases respectively. As the sampling operation is random, we opt not to deliberately pursue a uniform number of samples in each design. Notice that the above training and validation datasets have no intersection across designs. Refer to Section IV-A for more details.
Implementation Details. We train all models, including baseline methods, using Adam optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, the initial learning rate is set as 0.0008 and apply a low weight decay of $0.0001$. We gradually reduced the initial learning rate with the cosine annealing strategy. Datasets pre-processing follows [33]. Note that we do not impose any data augmentation transform during the training stage.
TABLE IV: Evaluation results on Mini-CircuitNet dataset. Red denotes percentage of improvement. $\downarrow$ means “lower is better”, $\uparrow$ means “higher is better”.
Method Mini-CircuitNet
@@
PowerNet [18] 0.171 0.62 0.75 0.70 0.60
PDNNet (Ours) 0.034(-80.1%) 0.87(+40.3%) 0.94(+25.3%) 0.75(+7.1%) 0.62(+3.3%)
MAVIREC [19] 0.056 0.82 0.92 0.71 0.58
PDNNet (Ours) 0.034(-39.3%) 0.87(+6.09%) 0.94(+2.2%) 0.75(+5.6%) 0.62(+6.9%)
@@ IV-B Main Experiment
We first compare our PDNNet with the state-of-the-art CNNs from the literature.  The first baseline for comparison is PowerNet [18], which is the first paper to perform dynamic IR drop prediction with deep-learning method.  The second is MAVIREC [19], a UNet-like model learning on dynamic IR drop with input vector, which is currently the state-of-the-art among the vector-based or vector-less methods. In general, the CircuitNet dataset expects neural networks to have strong generalization and representation capabilities due to its diverse design compositions. Thus, for evaluation metrics, we report both the accuracy metric (i.e., Normalized Mean Absolute Error (NMAE)), coefficient of determination (i.e., R-square ($R^{2}$)) and the graph evaluation criterion (i.e., Pearson, Spearman, Kendall correlation coefficient), to comprehensively assess the models’ performance. We abbreviate Pearson to Pear for shorten, Spearman and Kendall convert to Spea and Kend as the same. We calculate NMAE metric as $\frac{1}{N}\sum_{i=0}^{{N}-1}\frac{{\left|y_{i}-\tilde{y}_{i}\right|}}{max(y_{i})-min(y_{i})}$.
@@
Accuracy Comparison. The Table IV shows the results. Compared to the PowerNet, PDNNet achieves over 71% improvement (0.031 vs. 0.107) in NMAE metrics, which depicts the incomparable generalization of our PDNNet, other metrics also support this viewpoint. The weak performance of PowerNet is primarily due to its simplistic network design and non-mainstream 3D learning strategy, which limits its ability to learn deeper features related to IR drop. The typical CNN architecture, MAVIREC, still has a great margin compared with our PDNNet (0.038 vs. 0.031 in NMAE, 0.47 vs. 0.90 in $R^{2}$, 0.89 vs. 0.95 in Pearsonr, etc). It is worth noting that even in the experiments using Mini-CircuitNet with fewer samples (rightmost column of Table IV), our PDNNet has consistent and significant performance gains over all the five metrics, which implicitly demonstrates the limitation of overlooking PDN attributes during IR drop representation. Accuracy-computation comparison is illustrated in Figure 7. Visualization comparison results are shown in Figure 6. The figures from the left to right columns are macro positions in the floorplan, results from MAVIREC , results from our PDNNet, and ground truth label. The comparison results clearly show that our PDNNet generates IR drop prediction map with high structural fidelity, which demonstrates the effectiveness of our method.
@@
AUC Comparison. We also calculate the Area Under the Curve (AUC) score to measure model representation performance, PDNNet achieves 0.9748, outperforming the result of 0.9599 in MAVIREC.
Case 1
Refer to caption Refer to caption Refer to caption Refer to caption
Case 2
Refer to caption Refer to caption Refer to caption Refer to caption
Case 3
Refer to caption Refer to caption Refer to caption Refer to caption
Floorplan  MAVIREC [19] Our PDNNet   Ground Truth
Figure 6: Qualitative evaluation of CircuitNet dataset. The green polygons in the floorplan layout represent macros. Our PDNNet generates IR drop prediction map with high structural fidelity, which demonstrates the effectiveness of our method.
Accuracy-computation trade-off. Accuracy and computation cost are two crucial criteria for comprehensive model evaluation. The whole accuracy-computation trade-off of different methods is illustrated in Figure 7. We compare the accuracy and inference time between commercial tools, other compared methods, PDNNet variants, and our PDNNet. The inference time of our PDNNet is slightly higher than the MAVIREC (0.088s vs. 0.065s), while significantly lower than PowerNet and commercial tool Voltus (0.088s vs. 52.1s/60-120s). PDNNet and MAVIREC have similar inference costs as they both adopt 3D-CNN structures to aggregate temporal dimensions. For PowerNet, the unique max-CNN structure brings a large computation burden that costs much more inference time. The commercial tool, Voltus, uses a simulation-based method to acquire an IR drop distribution map and adopt it as our ground reference. Overall, the average runtime of PDNNet is around 0.11s (0.088s for model inference, 0.016s for data preprocessing), which performs over 545$\times$ speedup compared to commercial tool (1-2 minutes approximately).
@@ IV-C Ablation Studies
Figure 7: Accuracy-computation trade-off on Mini-CircuitNet for different methods in the inference phase. The horizontal axis shows the inference cost (i.e., whole inference time or model inference time) per sample. Whole inference time includes data processing and model inference.
Physical Intuitive GNN Branch Design (RQ1). We perform a preliminary exploration of GNN branch design by constructing two variants, a full layers of voltage-drop acquiring blocks and a full layers of neighbor influence blocks. We train the above models and PDNNet on Mini-CircuitNet with the same settings. Results in Table V, show that two variants have similar performance. The voltage-drop acquiring blocks is slightly higher than its neighbor influence counterpart (0.0348 vs. 0.0346) under the NMAE metric while exceeding in $R^{2}$ (0.8668 vs. 0.8665). However, by merging these distinctive layers in a specific manner, we observe a performance improvement in our PDNNet’s GNN branch (-2.9% and +0.3% in NMAE and $R^{2}$ respectively). These results reinforce the rationality of the formulation in Section II-A that shallow layers of GNN branch utilize current-related features to learn the preliminary representations of voltage drop, while deeper layers is to refine nodes’ voltage drop features influenced by their neighbors.
Dual-Branch Heterogeneous Architecture Analysis (RQ2). We perform a controlled study to assess PDNNet’s performance versus each variant model’s. The model set includes single-branch CNN and dual-branch CNN, plus single-branch GNN and dual-branch GNN. The dual-branch variants stand for using the same architecture as PDNNet while simply substituting with two identical branches (CNN or GNN). Table VI contains the comparison of the results, while the accuracy-speed comparison is shown in the inner box of  Figure 7.
TABLE V: Results of constructing GNN branch with different message-passing layers.
[…]
Full Voltage-Drop Acquiring Layers 0.0348 0.8668
Full Neighbor Influence Layers 0.0346 0.8665
Mixed GNN Layers (our GNN branch) 0.0336 0.8695
TABLE VI: Influence of GNN & CNN branch of PDNNet.
[…]
CNN 0.0447 0.8253
0.0361 0.8602
GNN 0.0420 0.8332
0.0347 0.8659
PDNNet 0.0336 0.8695
We have the following observations. First, although CNNs dominate the IR drop prediction in previous works, we interestingly observe that our GNN branch’s performance and inference speed both match or exceed its CNN-based counterpart, whilst with a simple PDN perception and relatively shallow layers (two GNN layers, each accompanied with one GNN block). This phenomenon is largely due to our specifically designed PDNGraph data structure which effectively mimics current flow and fine-grained cell-PDN relation within the PDN grid. Second, compared to a single-branch, dual-branch architecture indeed boosts the prediction performance (19.2% gains in CNN, 17.4% gains in GNN, under NMAE). However, both of these variants do not outperform our heterogeneous PDNNet (0.0361/0.0347 vs. 0.0336, and 0.8602/0.8659 vs. 0.8695), which demonstrate the PDN configuration and dynamic current pattern representations are two complementary aspects for effective IR drop prediction.
@@ V Conclusion
This paper introduces a novel approach to accurately predict dynamic IR drop on the power delivery network in IC design. We present a novel graph structure, PDNGraph, to properly represent the PDN configuration and the fine-grained cell-PDN relation. Moreover, we propose a novel network, PDNNet, a heterogeneous network with GNN-CNN branches to favorably perceive the above crucial features. Extensive experiments on a large public dataset demonstrate that our PDNNet outperforms the existing state-of-the-art CNN-based methods by a large margin (up to 39.3% in NMAE) and achieves a 545$\times$ speedup compared to the commercial tool with minimal error. We hope our work can foster further inspiration for EDA researchers.

### v2

@@ IV-A Experimental Setting
TABLE III: Evaluation results on CircuitNet dataset. Our PDNNet achieves consistent performance gain on five evaluation metrics in both large and small datasets. $\downarrow$ means “lower is better”, $\uparrow$ means “higher is better”.
Method CircuitNet [34]
@@
PowerNet [19] 0.149 0.32 11.60 0.56 0.77 0.83 0.70 0.503
MAVIREC [20] 0.039 0.81 18.27 0.68 0.91 0.85 0.77 0.942
MAVIREC+image-based PDN feature 0.037 0.82 18.55 0.70 0.91 0.86 0.78 0.946
PDNNet (Ours) 0.028 0.84 19.35 0.72 0.92 0.87 0.81 0.953
Datasets. CircuitNet [34, 35] is the large-scale public dataset in development that provides IC design used in the real industrial scenario. Unlike most previous work using small datasets for validation, CircuitNet-N28 (28nm version) provides more than 10,000 samples from 6 basic RTL designs (54 different synthesized netlists) with a wide variety of macros (3 types), frequencies (50/200/500 MHz), and back-end flow (variant in settings of utilization, macro placement, power mesh, filler insertion, etc), which is favorably appropriate to examine the actual performance of dynamic IR drop prediction models. We follow the same protocol as CircuitNet [34], to split the training and testing sets. That is, 4 designs (RISCY-a, RISCY-b, RISCY-FPU-a, RISCY-FPU-b) with 7078 samples are for training, and the rest 2 designs (zero-riscy-a and zero-riscy-b) with 3164 samples are for testing. CircuitNet-N28 leverages Innovus and Voltus flow to obtain vectorless dynamic IR drop solutions. In our experiment, we use the power feature that was initially provided in CirucitNet which incorporates 5 types of power patterns with 24 channels. Further, we extract the PDN features from the Design Exchange Format (DEF) files while incorporating the power information to complete our PDNGraph construction. We also use the whole dataset for conducting the ablation studies in Section IV-C. Notice that the above training and validation datasets have no intersection across designs. Refer to Section III-B1 for more details.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
    MAVIREC [20] Our PDNNet   Ground Truth
Figure 8: Qualitative evaluation of CircuitNet dataset. Our PDNNet generates IR drop prediction map with high structural fidelity, which demonstrates the effectiveness of our method.
TABLE IV: Results of constructing GNN branch with different message-passing layers.
[…]
Full Voltage-Drop Acquiring Layers 0.0286 0.8343 19.2704 0.7156 0.9173 0.8562 0.7902
Full Neighbor Influence Layers 0.0287 0.8316 19.2119 0.7151 0.9166 0.8492 0.7832
Mixed GNN Layers (our GNN branch) 0.0282 0.8369 19.3483 0.7190 0.9186 0.8709 0.8064
TABLE V: Influence of GNN & CNN branch of PDNNet.
[…]
CNN 0.0370 0.8100 18.3112 0.6956 0.9130 0.8646 0.7890
0.0358 0.8193 18.8091 0.7004 0.9122 0.8641 0.7904
GNN 0.0365 0.8120 18.7100 0.6897 0.9066 0.8865 0.7911
0.0336 0.8231 18.9977 0.6983 0.9119 0.8690 0.7816
PDNNet 0.0282 0.8369 19.3483 0.7190 0.9186 0.8709 0.8064
Implementation Details. We train our model, using Adam optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, the initial learning rate is set as 0.0008 and apply a low weight decay of $0.0001$. Other baseline methods are all trained to achieve convergence. We gradually reduced the initial learning rate with the cosine annealing strategy. Datasets pre-processing follows [34]. Note that we do not impose any data augmentation transform during the training stage. As for the training loss, we adopt L1 and Dice (Dice similarity coefficient)  [36] loss and treat them with equal weight. Note that our modified network outputs an IR drop map consistent in dimensions with the layout, enhancing its validity. For a fair comparison, all baseline methods are modified to adopt different size output.
@@ IV-B Main Experiment
We first compare our PDNNet with the state-of-the-art CNNs from the literature. The first baseline for comparison is PowerNet [19], which is the first paper to perform dynamic IR drop prediction with deep-learning method. The second is MAVIREC [20], a UNet-like model learning on dynamic IR drop with input vector, which is currently the state-of-the-art among the vector-based or vector-less methods. The last is incorporate PDN density map and power pad effective distance map from IREDGe[21] as additional features into the MAVIREC network. This method serves as a comparison with our GNN approach for perceiving the PDN structure. In general, the CircuitNet dataset expects neural networks to have strong generalization and representation capabilities due to its diverse design compositions. Thus, for evaluation metrics, we report both the accuracy metric (i.e., Normalized Mean Absolute Error (NMAE)), coefficient of determination (i.e., R-square ($R^{2}$)), image similarity (i.e., Peak Signal-to-Noise Ratio, Structural Similarity Index Measure), and the graph evaluation criterion (i.e., Pearson, Spearman, Kendall correlation coefficient), and area under the curve (AUC) to comprehensively assess the models’ performance. We abbreviate Pearson to Pear for shorten, Spearman and Kendall convert to Spea and Kend as the same. PSNR and SSIM are commonly used terms to refer to the above image similarity indexes. We calculate NMAE metric as $\frac{1}{N}\sum_{i=0}^{{N}-1}\frac{{\left|y_{i}-\tilde{y}_{i}\right|}}{max(y_{i})-min(y_{i})}$.
Accuracy Comparison.
@@
The Table III shows the results. Compared to the PowerNet, PDNNet achieves improvement (0.028 vs. 0.149) in NMAE metrics, which depicts the incomparable generalization of our PDNNet, other metrics also support this viewpoint. The weak performance of PowerNet is primarily due to its simplistic network design and non-mainstream 3D learning strategy, which limits its ability to learn deeper features related to IR drop. The typical CNN architecture, MAVIREC, still has a great margin compared with our PDNNet (0.039 vs. 0.028 in NMAE, 0.81 vs. 0.84 in $R^{2}$, 18.27dB vs. 19.35dB in PSNR, 0.68 vs. 0.72 in SSIM, 0.91 vs. 0.92 in Pearson, etc). When comparing with image-based PDN representation methods. PDNNet also present superiority in all metrics. It demonstrates the feature extraction process of GNNs for PDN structures can be viewed as a learnable representation, offering superior expressiveness compared to fixed image-based features. It is worth noting that our PDNNet has consistent and significant performance gains over all seven metrics, which demonstrates the complementary role of PDN attributes perception and dynamic current and during IR drop representation. Visualization comparison results are shown in Figure 8. The figures from the left to right columns are results from MAVIREC [20], results from our PDNNet, and ground truth label. The comparison results clearly show that our PDNNet generates IR drop prediction map with high structural fidelity, which demonstrates the effectiveness of our method.
AUC Comparison.
@@
We also calculate the AUC score to measure model representation performance, PDNNet achieves 0.953, outperforming the result of 0.946 in MAVIREC+image-based feature, 0.942 in MAVIREC, and 0.503 in PowerNet.
@@ IV-C Ablation Studies
Physical Intuitive GNN Branch Design (RQ1).
We perform a preliminary exploration of GNN branch design by constructing two variants, a full layer of voltage-drop acquiring blocks and a full layer of neighbor influence blocks. We train the above models and PDNNet on the whole CircuitNet with the same settings. Results in Table IV, show that two variants have similar performance. The branch with full neighbor influence layers is slightly higher than its voltage-drop acquiring counterpart in NMAE metric (0.0287 vs. 0.0286), while lower in $R^{2}$ (0.8316 vs. 0.8343). However, by merging these distinctive layers in a specific manner, we observe a performance improvement in our PDNNet’s GNN branch over all metrics. These results reinforce the rationality of the formulation in Section II-A that shallow layers of GNN branch utilize current-related features to learn the preliminary representations of voltage drop, while deeper layers are to refine nodes’ voltage drop features influenced by their neighbors.
Dual-Branch Heterogeneous Architecture Analysis (RQ2).
We perform a controlled study to assess PDNNet’s performance versus each variant model’s. The model set includes single-branch CNN and dual-branch CNN, plus single-branch GNN and dual-branch GNN. The dual-branch variants stand for using the same architecture as PDNNet while simply substituting with two identical branches (CNN or GNN). Table V contains the comparison of the results. We have the following observations. First, although CNNs dominate the IR drop prediction in previous works, we interestingly observe that our GNN branch’s performance in a single branch has similarly performance compared its CNN-based counterpart, whilst with a simple PDN perception and relatively shallow layers (two GNN layers, each accompanied by one GNN block. The parameter is also less than CNN, 41.15K vs. 17.07M). This phenomenon is largely due to our specifically designed PDNGraph data structure which effectively mimics current flow and fine-grained cell-PDN relation within the PDN grid. In dual-branch architecture, despite the NMAE, $R^{2}$ and PSNR, the dual-branch CNN have similar performance compared with GNN counterpart in other metrics.However, the parameter in dual-branch CNN is far more than GNN 34.13M vs. 78.02K. Second, compared to a single-branch, dual-branch architecture indeed boosts the prediction performance. However, both of these variants do not outperform our heterogeneous PDNNet (0.0358/0.0336 vs. 0.0282, and 0.8193/0.8231 vs. 0.8396) with 17.1M parameters, which demonstrate the PDN configuration and dynamic current pattern representations are two complementary aspects for effective IR drop prediction.
@@ V Conclusion
This paper introduces a novel approach to accurately predict dynamic IR drop on the power delivery network in IC design. We present a novel graph structure, PDNGraph, to properly represent the PDN configuration and the fine-grained cell-PDN relation. Moreover, we propose a novel network, PDNNet, a heterogeneous network with GNN-CNN branches to favorably perceive the above crucial features. Extensive experiments on a large public dataset demonstrate that our PDNNet outperforms the existing state-of-the-art CNN-based methods by a large margin and achieves a 545$\times$ speedup compared to the commercial tool with minimal error. We hope our work can foster further inspiration for EDA researchers.